# Credit Risk Meets Large Language Models: Building a Risk Indicator from Loan Descriptions in P2P Lending

### v1

@@ 7.1. Effect of including the BERT score
Table 13. Agreement and disagreement of the classifier with and without the BERT score.
Quant. + Categ. + BERT
Hits Misses
Quant. + Categ. Hits 63907 5004
Misses 8046 42144
Disagreement % 10.96%
The performance metrics in Table 12 indicate that incorporating the BERT score results in improved precision but diminished recall. These values suggest a discernible impact on the classifier’s behavior with the introduction of the new variable. Table 13 examines whether both classifiers yield the same outcome and whether it is accurate (hit) or inaccurate (miss). Notably, the cells in the inverse diagonal exhibit a significant number of instances, accounting for 11% of the total. When assessing the disparity in the accuracy of both classifiers, the inclusion of the BERT score leads to a modest 2% improvement. Therefore, while the incorporation of the BERT score enhances balanced performance, it does so not only by rectifying misses from the other classifier but also by influencing predictions for other instances.
Table 14 shows an additional experiment in which classifiers only use one kind of input variable. While the classifier using the quantitative variables is clearly the best, the classifier that uses just the BERT score obtains slightly better results than the one using the four categorical variables [^11] The difference in the AUC is significant at the 0.01 level according to the DeLong test.. This result is noteworthy given the well-known effectiveness and the meaningful nature of the qualitative variables.
Table 14. Performance considering only a kind of variables.
Metric Quant. Categ. BERT score
BACC 0.6062 0.5486 0.5490
AUC 0.6457 0.5656 0.5714
F1 0.3192 0.2759 0.2601
Precision 0.2138 0.1746 0.1877
Recall 0.6300 0.6563 0.5153
Accuracy 0.5896 0.4738 0.5724
7.2. Feature importance and explainability
To further analyze the impact of the BERT score, in this section, we analyze whether the models’ use of the input variables changes. Figure 4 shows the feature importance given by XGBoost both with and without the incorporation of the BERT score. Notably, the BERT score emerges as the third most influential variable, accounting for 6% importance. Moreover, the order and magnitude of importance for other variables exhibit shifts. Specifically, the borrower’s annual revenue diminishes in importance following the inclusion of the BERT score, while the significance of the ‘credit card’ purpose increases. This observation underscores the significant influence of the BERT score on the classifier.
@@ 7.2. Feature importance and explainability
Figure 4. Feature importances of the XGBoost classifiers.
To gain a deeper understanding of the role of variables, we employ SHAP values (Lundberg and Lee, 2017). Figure 5 provides a visualization of the SHAP values for the 10 variables with the most significant impact on the model’s output, comparing models with and without the inclusion of the BERT score.
@@
Figure 5. SHAP values of the XGBoost classifiers.
It is noteworthy that Figure 4(a) resembles a comparable figure presented in (Ariza-Garzón et al., 2020), wherein SHAP values were utilized in a lending model on Lending Club [^12] Our dataset is more constrained, only considering loans with textual descriptions, as explained in Section 4.. Examining Figure 4(b), we find that the BERT score ranks as the fourth variable in terms of impact. A closer inspection of the violin plots reveals that the incorporation of BERT does not substantially alter the way variables influence the target class.
Figure 6. Dependence plot between SHAP values and BERT score.
@@
7.3. Impact across the purpose categories
@@
Figure 7. BACC changes by purpose after including the BERT score.
Figure 7 illustrates the changes in balanced accuracy upon incorporating the BERT score for each loan purpose. Notably, the ‘other’ category shows a modest improvement of 2.11%. However, several other categories exhibit more substantial enhancements, including ‘educational’ (9.22%), ‘moving’ (5.84%), ‘medical’ (3.84%), and ‘small business’ (3.40%). These categories share a commonality—the more detailed specification of purpose or a deeper understanding of the borrower’s situation contributes to a more precise delineation of the default risk. For instance, in education loans, factors like the field of study and the educational institution may play a crucial role in determining employability and salaries, thereby enabling a more accurate assessment of the default risk. The same happens in the case of moving to a new place, having to pay medical expenses, or obtaining funds for a business.
@@
In each of these examples, the corresponding BERT scores fall below 0.4. As demonstrated in Figure 6, when the score is below this value, the classifier will have a stronger driver to categorize the loan as non-default. All the examples offer more details on the situation and the purpose of the loan. However, given the black-box nature of our BERT-based approach, the specific factors influencing the BERT score cannot be precisely determined. Furthermore, as we have shown, the score is also related to linguistic factors like tone, or language accuracy, and all the texts shown are accurate, readable, and written in a confident and objective tone. However, previous evidence in the literature offered mixed results about the influence of linguistic factors as default predictors (Gao and Lin, 2015; Wang et al., 2016; Dorfleitner et al., 2016; Siering, 2023).
8. Conclusions
In this paper, we have introduced a novel approach leveraging state-of-the-art natural language processing techniques to enhance credit risk models in P2P lending. Specifically, we employed BERT, a widely adopted LLM, to analyze the loan descriptions provided by borrowers. We successfully fine-tuned BERT to distinguish between defaulted and non-defaulted loans and used the model’s output as a risk score. Incorporating the BERT-based risk score improves the classification performance of a conventional granting model in a statistically significant manner. In fact, employing only this score as an input variable yields superior performance compared to considering the categorical variables alone. These results showcase the potential of LLMs to enhance lending models that rely on traditional variables.
Our approach can be easily applied without the need for manual annotation, which is a time-consuming process that also introduces annotator subjectivity. Once the LLM is trained, it instantly generates the risk score for a loan description. As we have illustrated, the score takes into account both linguistic factors and content aspects that relate to the creditworthiness of the borrower’s situation and purpose. However, the true drivers behind a given score or any potential biases remain uncertain. This lack of transparency and explainability is crucial not only for regulatory compliance but also for fostering trust among borrowers and lenders (ROFIEG, 2019).
@@ 8. Conclusions
These facts underscore the need for further exploration and integration of cutting-edge methodologies to enhance the effectiveness of risk-return assessments in the dynamic landscape of P2P lending or even in traditional financial institutions. Several directions deserve to be explored to refine the predictive capability of the model and the understanding of loan applicants’ conditions.
@@
A potential research avenue involves delving into more advanced LLMs. For instance, encoder-only models like RoBERTa (Liu et al., 2019) could improve the model’s capacity to capture more intricate linguistic patterns. Furthermore, the rise of generative AI has spurred interest in exploring decoder models for classification purposes. Although these models were not originally designed for such tasks, recent advancements have shown promising approximations. For example, CARP (Clue And Reasoning Prompting) (Sun et al., 2023) utilizes in-context learning, generating textual responses conditioned on a provided prompt with a few annotated examples (few-shot learning). Remarkably, this approach improved results in classification tasks without the need for transfer learning and fine-tuning and could be applied to risk scoring.
Another promising area of research is to use these language models in a less opaque way, such as through the use of topic models. While many topic models struggle to generate interpretable topics with large and heavily tailed vocabularies, novel approaches that combine traditional topic modeling and word embeddings (Dieng et al., 2020) can address these challenges. Applying such topic modeling approaches to loan descriptions could facilitate the identification of situations with varying risk levels. These identified topics can then be incorporated as additional variables in the classifier model, enhancing overall predictive performance without compromising the demanded transparency.

### v2

@@ 7.1 Analysis of the classification performance
The performance metrics in Table 12 indicate that incorporating the BERT score results in improved precision but diminished recall. However, attributing this change in precision-recall behavior solely to the BERT score requires careful consideration. This caution arises from our observations within our dataset, where classifiers with different hyperparameters and similar near-optimal balanced accuracy values have demonstrated varying precision-recall behaviors, suggesting that this may also be the case here.
Table 13 shows an additional experiment in which XGBoost classifiers are trained and optimized using only one kind of input variable. While the classifier using the quantitative variables is clearly the best, the classifier that uses just the BERT score obtains slightly better results than the one using the four categorical variables (the AUC difference is significant at the 0.01 level according to the DeLong test). This result is noteworthy given the well-known effectiveness and the meaningful nature of the qualitative variables.
Table 13 also shows the result of an XGBoost classifier using the textual features presented in Table 11, namely: polarity, subjectivity, word count, and readability score. This classifier is outperformed by the XGBoost that uses the BERT score in terms of balanced accuracy and AUC (significant at the 0.01 level according to the DeLong test). This finding underscores the superior ability of the fine-tuned LLM to leverage textual descriptions and extract relevant information for the classification task.
Table 13: Performance of the XGBoost considering only a subset of variables.
Metric Quant. Categ. Text. BERT score
BACC 0.6062 0.5486 0.5258 0.5490
AUC 0.6457 0.5656 0.5309 0.5714
F1 0.3192 0.2759 0.2534 0.2601
Precision 0.2138 0.1746 0.1665 0.1877
Recall 0.6300 0.6563 0.5302 0.5153
Accuracy 0.5896 0.4738 0.5227 0.5724
7.2 Feature importance and explainability
To further analyze the impact of the BERT score, in this section, we analyze whether the models’ use of the input variables changes. Figure 4 shows the feature importance given by XGBoost both with and without incorporating the BERT score. Notably, the BERT score emerges as the third most influential variable, accounting for 6% importance. Moreover, the order and magnitude of importance for other variables exhibit shifts. Specifically, the borrower’s annual revenue diminishes in importance following the inclusion of the BERT score, while the significance of the ‘credit card’ purpose increases. This observation underscores the significant influence of the BERT score on the classifier.
@@ 7.2 Feature importance and explainability
Figure 4: Feature importances of the XGBoost classifiers.
To gain a deeper understanding of the role of variables, we employ SHAP values [48]. Figure 5 provides a visualization of the SHAP values for the 10 variables with the most significant impact on the model’s output, comparing models with and without including the BERT score.
@@
Figure 5: SHAP values of the XGBoost classifiers.
It is noteworthy that Figure 5(a) resembles a comparable figure presented in [41], wherein SHAP values were utilized in a lending model on Lending Club [^11] Our dataset is more constrained, only considering loans with textual descriptions, as explained in Section 4.. Examining Figure 5(b), we find that the BERT score ranks as the fourth variable in terms of impact. A closer inspection of the violin plots reveals that including BERT does not substantially alter how variables influence the target class.
Figure 6: Dependence plot between SHAP values and BERT score.
@@
7.3 Impact across the purpose categories
@@
Figure 7: BACC changes by purpose after including the BERT score.
Figure 7 illustrates the changes in balanced accuracy upon incorporating the BERT score for each loan purpose. Notably, the ‘other’ category shows a modest improvement of 2.11%. However, several other categories exhibit more substantial enhancements, including ‘educational’ (9.22%), ‘moving’ (5.84%), ‘medical’ (3.84%), and ‘small business’ (3.40%). Given the black-box nature of our model, it is difficult to ascertain the reasons why these categories have improved more than the others. However, we posit that these categories share a commonality—the more detailed specification of purpose or a deeper understanding of the borrower’s situation contributes to a more precise delineation of the default risk. For instance, in education loans, factors like the field of study and the educational institution may play a crucial role in determining employability and salaries, thereby enabling a more accurate assessment of the default risk. The same happens in the case of moving to a new place, having to pay medical expenses, or obtaining funds for a business.
@@
In each of these examples, the corresponding BERT scores fall below 0.4. As demonstrated in Figure 6, the classifier will have a stronger driver to categorize the loan as non-default when the score is below this value. As all the examples offer more details on the situation and the purpose of the loan, their further clarity could explain their low BERT score. In addition, all the texts shown are accurate, readable, and written in a confident and objective tone. While previous evidence in the literature offered mixed results about the influence of linguistic factors as default predictors [3, 18, 4, 7], we have shown that our BERT score is significantly related to them. While these explanations are sound, given the black-box nature of our BERT-based approach, the explanations influencing the BERT score of each text cannot be precisely determined. This aspect needs to be improved to comply with the requirements of the regulators and foster trust among the end-users.
8 Conclusions
In this paper, we have introduced a novel approach that highlights the value of state-of-the-art natural language processing techniques in enhancing credit risk models. Our approach fine-tunes BERT using loan descriptions to generate a risk score that effectively discriminates between defaulted and non-defaulted loans. Furthermore, we show that integrating the BERT-based risk score with traditional variables significantly enhances the performance of conventional granting models. These results underscore the potential of LLMs to augment lending models, emphasizing their invaluable role in harnessing the richness of textual data to improve credit risk assessment.
Our approach can be easily applied without manual annotation, which is a complex task that also introduces annotator subjectivity. As our analysis shows, the score incorporates both linguistic factors and content aspects that relate to the creditworthiness of the borrower’s situation and purpose. Additionally, while fine-tuning the LLM is a time-consuming process that requires GPU processing, generating predictions, such as the risk score for a loan description, can be done instantly with standard home equipment.
@@ 8 Conclusions
Our work represents just the first steps in this endeavor and underscores the necessity for further exploration and integration of cutting-edge methodologies to enhance the effectiveness of risk-return assessments in the dynamic landscape of P2P lending or even in traditional financial institutions. Several directions deserve to be explored to refine the predictive capability of the model and the understanding of loan applicants’ conditions.
An important limitation of our work is the lack of transparency regarding the factors contributing to a given risk score and its potential biases. These issues hinder its application in real-world contexts, where a comprehensive understanding of the score generation process and assurance that it does not introduce biases related to gender, ethnicity, financial inclusion, or other factors are essential. Thus, efforts to enhance the transparency and explainability of the risk score are crucial not only for regulatory compliance but also for fostering trust among borrowers and lenders [49].
At this respect, LLM-based topic modeling tools such as BERTopic [50] could be used to gain an understanding of the risk scores. For example, topic modeling could identify whether a loan description is related to “risky topics” that in turn could be incorporated as input variables in the granting model. Similarly, the embeddings from BERT or other LLMs could be used to generate interpretable topics from large, complex vocabularies, as other works have done [51]. Applying such topic modeling approaches to loan descriptions could facilitate the identification of situations with varying risk levels. These topics could then serve as additional variables in classification models, improving predictive performance while maintaining the transparency required in financial assessments.
@@
Another potential research avenue involves delving into more advanced LLMs. For instance, encoder-only models like RoBERTa [37] could improve the model’s capacity to capture more intricate linguistic patterns. Furthermore, the rise of generative AI has spurred interest in exploring decoder models for classification purposes. Although these models were not originally designed for such tasks, recent advancements have shown promising approximations. For example, CARP (Clue And Reasoning Prompting) [52] utilizes in-context learning, generating textual responses conditioned on a provided prompt with a few annotated examples (few-shot learning). Remarkably, this approach improved results in classification tasks without the need for transfer learning and fine-tuning and could be applied to risk scoring.
Future work should also address the economic implications of our findings by integrating cost- or profit-sensitive approaches [53, 42]. These methods allow for the inclusion of the financial impacts of classification outcomes. Investigating whether improvements in model performance by including textual descriptions translate into better financial outcomes would enhance the real-world applicability of our findings.